# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ 1.5 Optimization of parameters
$=\log E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\left[\frac{p(\mathbf{x}_{0:T},\boldsymbol{\theta}_{1:T}|\mathbf{w}^{*})}{p(\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w})}\right]$
@@
$=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log p(\mathbf{x}_{0:T},\boldsymbol{\theta}_{1:T}|\mathbf{w}^{*})-E_{\boldsymbol{\theta}_{1:T}|\mathbf{w}}\log p(\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w})$ (128)
The first term is called Q-function:
${\tilde{Q}(\mathbf{Q})}$ $=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log[p(\mathbf{x}_{0:T},\boldsymbol{\theta}_{1:T}|\mathbf{Q})]$
$=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log[p(\mathbf{x}_{0:T}|\boldsymbol{\theta}_{1:T},\mathbf{Q})]+E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log[p(\boldsymbol{\theta}_{1:T}|\mathbf{Q})]$ (129)
The second term is the entropy of the posterior density which is fixed at M-step. We thus optimize the hyperparameters that maximizes the Q-function. Hereafter, we denote the parameter $\mathbf{w}^{*}$ by $\mathbf{Q}$.
First, we focus on the first term related to the observation model. However, since the parameter $\mathbf{Q}$ does not influence the estimation of $\mathbf{x}_{1:T}$, this term may be neglected in the optimization.
${\tilde{Q}(\mathbf{Q})}$ $=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log[p(\mathbf{x}_{0:T},\boldsymbol{\theta}_{1:T}|\mathbf{Q})]$
$=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log[p(\mathbf{x}_{0:T}|\boldsymbol{\theta}_{1:T})]+E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{1:T},\mathbf{w}}\log[p(\boldsymbol{\theta}_{1:T}|\mathbf{Q})]$ (130)
@@
$\log p(\mathbf{x}_{0:T}|\boldsymbol{\theta}_{1:T})$ $=\log[\prod_{t=1}^{T}p(\mathbf{x}_{t}|\mathbf{x}_{t-1},\boldsymbol{\theta}_{T})]$
$=\log\prod_{t=1}^{T}\prod_{i=1}^{N}\exp[\sum_{l=1}^{L}\{(\boldsymbol{\theta}^{i}_{t})^{T}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\sum_{l=1}^{L}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\}]$ (131)
we derive the state model $p(\boldsymbol{\theta}_{t}|\boldsymbol{\theta}_{t-1},\mathbf{Q})$, which is defined as follows:
$p(\boldsymbol{\theta}_{1:T})=p(\boldsymbol{\theta}_{1}|\boldsymbol{\mu},\mathbf{\Sigma})\prod_{t=2}^{T}p(\boldsymbol{\theta}_{T}|\boldsymbol{\theta}_{T-1})$ (132)
$p(\boldsymbol{\theta}_{t}|\boldsymbol{\theta}_{t-1})=\prod_{i=1}^{N}p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1})=\prod_{i=1}^{N}\mathcal{N}(\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})$ (133)
With this,
@@
$\log[p(\boldsymbol{\theta}_{1:T}|\mathbf{Q})]$ $=\log\left[p(\boldsymbol{\theta}_{1}|\boldsymbol{\mu},\mathbf{\Sigma})\prod_{t=2}^{T}p(\boldsymbol{\theta}_{t}|\boldsymbol{\theta}_{t-1})\right]$
$=\log\left[\prod_{i=1}^{N}p(\boldsymbol{\theta}^{i}_{1}|\boldsymbol{\mu}^{i},\mathbf{\Sigma}^{i})\right]+\log\left[\prod_{t=2}^{T}\prod_{i=1}^{N}\mathcal{N}(\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})\right]$
$=\sum_{i=1}^{N}\log\mathcal{N}(\boldsymbol{\mu}^{i},\mathbf{\Sigma}^{i})+\sum_{t=2}^{T}\sum_{i=1}^{N}\log\left[\frac{1}{\sqrt{|2\pi\mathbf{Q}^{i}|}}\right]$
$\quad-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})$ (134)
We define the function $\tilde{Q}(\mathbf{Q})$} as follows:
$\tilde{Q}(\mathbf{Q})$ $=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\sum_{t=1}^{T}\sum_{i=1}^{N}\sum_{l=1}^{L}[(\boldsymbol{\theta}^{i}_{t})^{T}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\psi(\mathbf{x}^{l}_{t-1})]$
$+E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\sum_{t=2}^{T}\sum_{i=1}^{N}[-\frac{1}{2}\log|2\pi\mathbf{Q}^{i}|-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})]$
$+E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}[\sum_{i=1}^{N}-\frac{1}{2}\log|2\pi\mathbf{\Sigma}^{i}|-\frac{1}{2}(\boldsymbol{\theta}^{i}_{1}-\boldsymbol{\mu}^{i})^{\top}(\mathbf{\Sigma}^{i})^{-1}(\boldsymbol{\theta}^{i}_{1}-\boldsymbol{\mu}^{i})]$ (135)
$\frac{\partial\log|2\pi\mathbf{Q}^{i}|}{\partial\mathbf{Q}^{i}}=\frac{1}{|\mathbf{Q}^{i}|}\frac{\partial|\mathbf{Q}^{i}|}{\partial\mathbf{Q}^{i}}=\frac{1}{|\mathbf{Q}^{i}|}|\mathbf{Q}^{i}|(\mathbf{Q}^{i})^{\mathrm{-1}}=[(\mathbf{Q}^{i})^{-1}]^{\top}$ (136)
@@
$\frac{\partial}{\partial\mathbf{Q}^{i}}$ $(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})$
$=\frac{\partial(\mathbf{Q}^{i})^{-1}}{\partial\mathbf{Q}^{i}}\frac{\partial}{\partial(\mathbf{Q}^{i})^{-1}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})$
$=-\mathbf{Q}^{i^{-2}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top},$ (137)
we obtain
$\frac{\partial\tilde{Q}(\mathbf{Q})}{\partial\mathbf{Q}^{i}}=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{1:T},\mathbf{w}}\sum_{t=2}^{T}\left[-\frac{1}{2}\left((\mathbf{Q}^{i})^{-1}\right)^{\top}+\frac{1}{2}\mathbf{Q}^{i^{-2}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}\right].$ (138)
@@
$\mathbf{Q}^{i}=\frac{1}{T-1}\sum_{t=2}^{T}E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{1:T},\mathbf{w}}\left[(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}\right]$ (139)
Next, we examine the calculation of the expected values.
$E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{1:T},\mathbf{w}}\left[(\boldsymbol{\theta}^{i}_{t})(\boldsymbol{\theta}^{i}_{t})^{\top}-(\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t})^{\top}-(\boldsymbol{\theta}^{i}_{t})(\boldsymbol{\theta}^{i}_{t-1})^{\top}+(\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t-1})^{\top}\right]$ (140)
Using the following definitions of the equal-time covariance matrix:
$\mathbf{W}^{i}_{t|T}=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\left[(\boldsymbol{\theta}^{i}_{t})(\boldsymbol{\theta}^{i}_{t})^{\top}\right]-(\boldsymbol{\theta}^{i}_{t|T})(\boldsymbol{\theta}^{i}_{t|T})^{\top},$ (141)
and the delayed covariance:
$\mathbf{W}^{i}_{t,t-1|T}$ $=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\left[(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|T})(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|T})^{\top}\right]$
$=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\left[(\boldsymbol{\theta}^{i}_{t})(\boldsymbol{\theta}^{i}_{t-1})^{\top}\right]-(\boldsymbol{\theta}^{i}_{t|T})(\boldsymbol{\theta}^{i}_{t-1|T})^{\top}.$ (142)
The optimal $\mathbf{Q}^{i}$ is written as
$\mathbf{Q}^{i}=\frac{1}{T-1}$ $\sum_{t=2}^{T}\left((\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})^{\top}\right$.
$\quad\left.+\mathbf{W}^{i}_{t|T}-\mathbf{W}^{i}_{t-1,t|T}-\mathbf{W}^{i}_{t,t-1|T}+\mathbf{W}^{i}_{t-1|T}\right),$ (143)
where $\mathbf{W}^{i}_{t-1,t|T}=(\mathbf{W}^{i}_{t,t-1|T})^{\top}$ . We compute the delayed covariance matrix following the method of De Jong and Mackinnon [95]:
$\mathbf{W}^{i}_{t,t-1|T}=\mathbf{W}^{i}_{t|t}(\mathbf{W}^{i}_{t+1|t})^{-1}\mathbf{W}^{i}_{t|T}$ (144)
This is the $\mathbf{Q}$ (or equivalently, $\mathbf{w}^{*}$) that maximizes the likelihood function $\tilde{Q}(\mathbf{Q})$}.
Similarly, we update $\boldsymbol{\Sigma}^{i}$ according to
1.6 Evaluation of log marginal likelihood functions
The convergence of the EM algorithm is evaluated using the log marginal likelihood. Consequently, an approximate solution for the log marginal likelihood of the kinetic Ising model is obtained. First, let the following marginal likelihood function $p(\mathbf{x}_{0:T}|\mathbf{w})$ be expressed as the mean and variance of the filter distribution.
$p(\mathbf{x}_{0:T}|\mathbf{w})$ $=p(\mathbf{x}_{0}|\mathbf{w})\prod_{t=2}^{T}p(\mathbf{x}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$
$=p(\mathbf{x}_{0}|\mathbf{w})\prod_{t=2}^{T}\int d\boldsymbol{\theta}_{t}p(\mathbf{x}_{t}|\mathbf{x}_{0:t-1},\boldsymbol{\theta}_{t},\mathbf{w})p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$
$=p(\mathbf{x}_{0}|\mathbf{w})\prod_{t=2}^{T}\int d\boldsymbol{\theta}_{t}p(\mathbf{x}_{t}|\mathbf{x}_{t-1},\boldsymbol{\theta}_{t})p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$
$=p(\mathbf{x}_{0}|\mathbf{w})\prod_{t=2}^{T}\prod_{i=1}^{N}\int d\boldsymbol{\theta}^{i}_{t}\prod_{l=1}^{L}p({x}^{l}_{i,t}|\mathbf{x}_{0:t-1},\boldsymbol{\theta}^{i}_{t},\mathbf{w})p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w}),$ (146)
Note that in the integrand, if $\boldsymbol{\theta}^{i}_{t}$ is known, the data $\mathbf{x}_{0:t-2}$ and the parameters $\mathbf{w}$ do not affect the estimation of $x^{l}_{i,t}$. Hence, these terms are omitted, and we have
$p({x}^{l}_{i,t}|\mathbf{x}_{0:t-1},\boldsymbol{\theta}^{i}_{t},\mathbf{w})=p({x}^{l}_{i,t}|\mathbf{x}^{l}_{t-1},\boldsymbol{\theta}^{i}_{t})$ (147)
The observed model and the one-period prediction distribution in the integral function can be summarized as
$\prod_{l=1}^{L}p({x}^{l}_{i,t}|\mathbf{x}^{l}_{t-1},\boldsymbol{\theta}^{i}_{t})=$ $\prod_{l=1}^{L}\exp[\theta_{i,t}x^{l}_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x^{l}_{it}x^{l}_{j,t-1}-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})]$
$=\exp[(\boldsymbol{\theta}^{i}_{t})^{T}\sum_{l=1}^{L}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\sum_{l=1}^{L}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})]$ (148)
$p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$ $=\frac{1}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}\exp[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})]$ (149)
Substituting (1.6) and (LABEL:onstep_m_in_L) into (1.6), we obtain
$p(\mathbf{x}_{0:T}|\mathbf{w})$ $=p(\mathbf{x}_{0}|\mathbf{w})\int\prod_{i=1}^{N}\prod_{t=2}^{T}d\boldsymbol{\theta}^{i}_{t}\frac{1}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}\exp[(\boldsymbol{\theta}^{i}_{t})^{T}\sum_{l}^{L}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})$
$-\sum_{l}^{L}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})]$ (151)
Now, we define the function $q$ as follows.
$q(\boldsymbol{\theta}^{i}_{t})$ $=(\boldsymbol{\theta}^{i}_{t})^{T}\sum_{l=1}^{L}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\sum_{l=1}^{L}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$
$\quad-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})$ (152)

### v3

@@ 1.5 Optimization of hyperparameters
$=\log E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\frac{p(\mathbf{x}_{0:T},\boldsymbol{\theta}_{1:T}|\mathbf{w}^{*})}{p(\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w})}$
@@
$=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log p(\mathbf{x}_{0:T},\boldsymbol{\theta}_{1:T}|\mathbf{w}^{*})-E_{\boldsymbol{\theta}_{1:T}|\mathbf{w}}\log p(\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}).$ (S1.53)
The first term is called the Q-function:
${\tilde{Q}(\mathbf{w})}$ $=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log p(\mathbf{x}_{0:T},\boldsymbol{\theta}_{1:T}|\mathbf{Q})$
$=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log p(\mathbf{x}_{0:T}|\boldsymbol{\theta}_{1:T},\mathbf{Q})+E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\log p(\boldsymbol{\theta}_{1:T}|\mathbf{Q}).$ (S1.54)
The second term is the entropy of the posterior density, which is fixed at M-step. We thus optimize the hyperparameters that maximize the Q-function. More explicitly, the Q-function can be written as
$\tilde{Q}(\mathbf{w})$ $=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\sum_{t=1}^{T}\sum_{i=1}^{N}\sum_{l=1}^{L}[(\boldsymbol{\theta}^{i}_{t})^{T}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\psi(\mathbf{x}^{l}_{t-1})]$
$+E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\sum_{i=1}^{N}\left[-\frac{1}{2}\log|2\pi\boldsymbol{\Sigma}^{i}|-\frac{1}{2}(\boldsymbol{\theta}^{i}_{1}-\boldsymbol{\mu}^{i})^{\top}(\boldsymbol{\Sigma}^{i})^{-1}(\boldsymbol{\theta}^{i}_{1}-\boldsymbol{\mu}^{i})\right]$
$+E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\sum_{t=2}^{T}\sum_{i=1}^{N}\left[-\frac{1}{2}\log|2\pi\mathbf{Q}^{i}|-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})\right].$ (S1.55)
$\frac{\partial\log|2\pi\mathbf{Q}^{i}|}{\partial\mathbf{Q}^{i}}=\frac{1}{|\mathbf{Q}^{i}|}\frac{\partial|\mathbf{Q}^{i}|}{\partial\mathbf{Q}^{i}}=\frac{1}{|\mathbf{Q}^{i}|}|\mathbf{Q}^{i}|(\mathbf{Q}^{i})^{\mathrm{-1}}=(\mathbf{Q}^{i})^{-1},$ (S1.56)
and
$\frac{\partial}{\partial\mathbf{Q}^{i}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})$ $=\frac{\partial(\mathbf{Q}^{i})^{-1}}{\partial\mathbf{Q}^{i}}\frac{\partial}{\partial(\mathbf{Q}^{i})^{-1}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})$
$=-(\mathbf{Q}^{i})^{-2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top},$ (S1.57)
@@
$\frac{\partial\tilde{Q}(\mathbf{w})}{\partial\mathbf{Q}^{i}}=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{1:T},\mathbf{w}}\sum_{t=2}^{T}\left[-\frac{1}{2}(\mathbf{Q}^{i})^{-1}+\frac{1}{2}(\mathbf{Q}^{i})^{-2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}\right].$ (S1.58)
$\mathbf{Q}^{i}=\frac{1}{T-1}\sum_{t=2}^{T}E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{1:T},\mathbf{w}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}.$ (S1.59)
We note that the expectation in the above equation can be decomposed into
$E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{1:T},\mathbf{w}}\left[\boldsymbol{\theta}^{i}_{t}(\boldsymbol{\theta}^{i}_{t})^{\top}-\boldsymbol{\theta}^{i}_{t-1}(\boldsymbol{\theta}^{i}_{t})^{\top}-\boldsymbol{\theta}^{i}_{t}(\boldsymbol{\theta}^{i}_{t-1})^{\top}+\boldsymbol{\theta}^{i}_{t-1}(\boldsymbol{\theta}^{i}_{t-1})^{\top}\right].$ (S1.60)
Hence, using the following definitions of the equal-time covariance matrix:
$\mathbf{W}^{i}_{t|T}=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\boldsymbol{\theta}^{i}_{t}(\boldsymbol{\theta}^{i}_{t})^{\top}-\boldsymbol{\theta}^{i}_{t|T}(\boldsymbol{\theta}^{i}_{t|T})^{\top},$ (S1.61)
and the delayed covariance:
$\mathbf{W}^{i}_{t,t-1|T}$ $=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|T})(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|T})^{\top}$
$=E_{\boldsymbol{\theta}_{1:T}|\mathbf{x}_{0:T},\mathbf{w}}\boldsymbol{\theta}^{i}_{t}(\boldsymbol{\theta}^{i}_{t-1})^{\top}-\boldsymbol{\theta}^{i}_{t|T}(\boldsymbol{\theta}^{i}_{t-1|T})^{\top},$ (S1.62)
the optimal $\mathbf{Q}^{i}$ is obtained as
$\mathbf{Q}^{i}=\frac{1}{T-1}$ $\sum_{t=2}^{T}\left[(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|T})^{\top}+\mathbf{W}^{i}_{t|T}-\mathbf{W}^{i}_{t-1,t|T}-\mathbf{W}^{i}_{t,t-1|T}+\mathbf{W}^{i}_{t-1|T}\right],$ (S1.63)
where $\mathbf{W}^{i}_{t-1,t|T}=(\mathbf{W}^{i}_{t,t-1|T})^{\top}$. We compute the lag-one smoothed covariance following the method of De Jong and Mackinnon [2]:
$\mathbf{W}^{i}_{t,t-1|T}=\mathbf{W}^{i}_{t|t}(\mathbf{W}^{i}_{t+1|t})^{-1}\mathbf{W}^{i}_{t|T}.$ (S1.64)
Similarly, we update $\boldsymbol{\Sigma}^{i}$ according to
$\boldsymbol{\Sigma}^{i}=\mathbf{W}^{i}_{1|T}+(\boldsymbol{\theta}^{i}_{1|T}-\boldsymbol{\mu})(\boldsymbol{\theta}^{i}_{1|T}-\boldsymbol{\mu})^{\top}.$ (S1.65)
1.6 Approximate log marginal likelihood function
The convergence of the EM algorithm was assessed using the log marginal likelihood. Below, we derive the approximate solution for the log marginal likelihood of the kinetic Ising model.
First, we note that the marginal likelihood function $p(\mathbf{x}_{0:T}|\mathbf{w})$ can be expressed as follows:
$=p(\mathbf{x}_{0})\prod_{t=1}^{T}\int d\boldsymbol{\theta}_{t}p(\mathbf{x}_{t}|\mathbf{x}_{0:t-1},\boldsymbol{\theta}_{t},\mathbf{w})p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$
$=p(\mathbf{x}_{0})\prod_{t=1}^{T}\int d\boldsymbol{\theta}_{t}p(\mathbf{x}_{t}|\mathbf{x}_{t-1},\boldsymbol{\theta}_{t})p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$
$=p(\mathbf{x}_{0})\prod_{t=1}^{T}\prod_{i=1}^{N}\int d\boldsymbol{\theta}^{i}_{t}\prod_{l=1}^{L}p({x}^{l}_{i,t}|\mathbf{x}_{t-1},\boldsymbol{\theta}^{i}_{t})p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w}).$ (S1.66)
The observation model and the one-step prediction density in the equation above are written as
$\prod_{l=1}^{L}p({x}^{l}_{i,t}|\mathbf{x}^{l}_{t-1},\boldsymbol{\theta}^{i}_{t})$ $=\prod_{l=1}^{L}\exp\left[\theta_{i,t}x^{l}_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x^{l}_{it}x^{l}_{j,t-1}-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right]$
$=\exp\left[(\boldsymbol{\theta}^{i}_{t})^{T}\sum_{l=1}^{L}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\sum_{l=1}^{L}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right],$ (S1.67)
@@ 1.6 Approximate log marginal likelihood function
$p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})$ $=\frac{1}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}\exp\left[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})\right].$ (S1.68)
Substituting Eqs.S1.67 and S1.68 into Eq.1.6, we obtain
$p(\mathbf{x}_{0:T}|\mathbf{w})$ $=p(\mathbf{x}_{0})\prod_{i=1}^{N}\prod_{t=1}^{T}\int d\boldsymbol{\theta}^{i}_{t}\frac{1}{\sqrt{|2\pi\mathbf{W}^{i}_{t|t-1}|}}$
@@
$\phantom{=}\cdot\exp\left[(\boldsymbol{\theta}^{i}_{t})^{T}\sum_{l=1}^{L}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\sum_{l=1}^{L}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})\right].$ (S1.69)
We now define the function $q(\boldsymbol{\theta}^{i}_{t})$ as follows:
$q(\boldsymbol{\theta}^{i}_{t})$ $=(\boldsymbol{\theta}^{i}_{t})^{T}\sum_{l=1}^{L}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\sum_{l=1}^{L}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1}).$ (S1.70)